please help with understanding fraction part of question...make an equation....(-3,4) (6,1) with slope of 1/3
Answer by stanbon(75887)

make an equation....(-3,4) (6,1) with slope of 1/3
-------------------
Comment: Those two points do not have a slope of 1/3
Their slope = (4-1)/(-3-6) = 3/-9 = -1/3
---------------------------------------------
Using that slope and one of the points and the form y = mx + b
you get:
4 = (-1/3)*-3 + b ; solve for "b":
4 = 1 + b
b = 3
------------------
Equation:
y = (-1/3)x + 3
========================
Cheers,
Stan H.